# Alpine: An adaptive language-agnostic pruning method for language models for code

### v1

@@ IV Experiments
Naturally, pruning tokens would result in some information loss. Through this research question, we aim to measure the extent of performance drop, in terms of accuracy, for example, that might occur, if any. -
@@
The proposed approach allows to choose whether tokens that are pruned to be partially kept by merging their representation into a single row or to be entirely removed from the sequence. In this research question, we investigate the impact of such a design choice. -
@@
Finally, an important reason behind model simplification is to allow computationally demanding models to run on relatively lesser-capable gpus (e.g., consumer-grade gpus) compared to those used in high-performance computing clusters (e.g., NVIDIA V or A). This exploration would ensure wider accessibility and would make the usage of these models more practical.
IV-A Model selection
The key selection criterion for the models is that they should be encoder-only Transformer-based language models for code that is publicly available. Based on the criteria, we identify the following models for our experiments.
CodeBERT [38]: It is a pre-trained language model that has the same architecture and a similar pretraining strategy as RoBERTa [43]. It comprises twelve Transformer layers, each containing a 12-headed mha layer and has a hidden dimension of . It was trained on bimodal data consisting of pairs of natural language and programming language across six programming languages from the CodeSearchNet dataset [44].
GraphCodeBERT [45]: GraphCodeBert extends CodeBert by including the data flow information within the source code input. It also includes an additional pre-training task where the objective is to predict whether a data flow edge exists between two nodes.
UniXCoder [31]: UniXCoder leverages a multi-modal input consisting of natural language, source code, and the flattened sequence of the abstract syntax tree of the code snippet. It is similar to the two other models that share the same internal architecture. The difference here (aside from the pre-training objectives and input representation during pre-training) is that UniXCoder allows for a larger context length, compared to for (Graph)CodeBERT.
IV-B Software Engineering Tasks
In this study, we choose two substantially different tasks from the CodeXGLUE benchmark [41] for our experiments.
Code clone detection: This task is taken from the filtered version of the BigCloneBenchmark dataset [46] released by Wang et al. [47]. This binary classification problem aims to predict whether two Java code snippets are clones. The dataset is composed of training pairs and pairs for testing. We randomly take a subsample of pairs for training and for evaluation; such random sampling is quite common in this domain [20, 48]. We use the same subsamples for all experiments during training and evaluation.
Defect prediction: Zhou et al. [39] proposed this dataset from a set of four large-scale open-source C repositories. The authors manually annotated the code snippets and covered multiple types of defects such as resource leaks and use-after-free vulnerabilities. We use the default pre-split sets of samples for training and for validation and testing.
@@ IV-B Software Engineering Tasks
IV-C Evaluation Metrics
To assess the predictive performance of the models on the aforementioned tasks, we use the same evaluation metrics that were reported by Lu et al. [41] in the CodeXGLUE benchmark. Specifically, we calculate the accuracy for the Devign dataset and F1-score for the BigCloneBenchmark. As for the computational efficiency, we report the number of floating points operations (flops) and Throughput. As stated in Section II-C the number flops quantifies the computational complexity. A model with a higher flop count requires more computing resources. Our goal here is to reduce such a count while maintaining the predictive performance as much as possible. The throughput refers to the number of input samples a model can process in a second. This metric is especially relevant during inference and model deployment.
IV-D Experimental Setting
@@
V Results
V-A RQ1: Computational Efficiency
@@
Impact on FLOPs Count: Table II(b) highlights the computational speedup offered by Alpine across the three models on the SE tasks. Performing pruning across all layers allows for the most computational savings. On average555Given the values are ratios, we report the geometric mean., it reduces the flops count by on the Devign dataset and and BigCloneBenchmark compared the baseline models (i.e., without any pruning). Specifically, CodeBert and GraphCodeBert exhibit roughly the same gain, where these models consume roughly half flops. UniXCoder also show significant gain, though slightly smaller ( and ), in terms of flops.
In Figure 3, we plot the average sequence length across all the twelve layers during inference. When pruning across all layers on the Devign dataset, the sequence length plot exhibits an exponential-like decay on CodeBert and GraphCodeBert. On the other hand, this reduction follows roughly a linear trend on BigCloneBenchmark across all models. What is interesting is that at the final layer, the sequences are quite compressed. Specifically, the highest average sequence length at the last layer () is obtained when using GraphCodeBert on the defect prediction task which is less than the none pruned average length of . In addition, aligned with the results of Table II(b), we can see that pruning at even-indexed layers yields shorter sequences than odd-indexed ones. Across all task-model combinations and all layers (except for - layer), Alpine produces more compressed sequences when applied at these layers compared to their odd-indexed counterpart.
Impact on Memory Footprint: To further evaluate Alpine’s impact from a computational efficiency standpoint, we plot the gpu memory consumption. Due to space constraints, we only report the figures regarding pruning across all layers.
V-B RQ2: Impact on Performance
In Table II(b), we report the predictive performance of the pruned models. On the Devign dataset, performing pruning at odd-indexed layers leads to a slight improvement when using CodeBert, with a % increase in accuracy compared to that of the non-pruned version. On the other hand, adding the pruning layer at even layers is more efficient for GraphCodeBert and UniXCoder, where they were able to maintain % and % of the original accuracy scores, respectively.
In the clone detection task, pruning at even-indexed layers achieves the highest F1-scores compared to other pruning strategies. It even managed to improve the clone detection performance by increasing the F1-scores of GraphCodeBert and UniXCoder by % and %.
V-C RQ3: Effect of Merging Prunable Tokens on Model Accuracy
During pruning, the representation of tokens marked to be pruned can be entirely removed or merged into one vector. To study the impact of token merging, we conduct experiments to measure its impact on the two tasks and report the results in Table III.
@@ V-C RQ3: Effect of Merging Prunable Tokens on Model Accuracy
V-D RQ4: Impact of Computing Environment
@@
Building upon these results, we further investigate the implications of the reduced time from an environmental perspective. In Table IV, we report the CO2 emission rates for both gpus with and without Alpine. The measurements were taken using the tool provided by Lacoste et al. [18].
@@
VI Threats to Validity
@@
To mitigate a potential threat, we used the same hyperparameters across all models, as reported in previous works [41], and utilized the same data splits for fine-tuning and testing.
@@
External threats to validity are concerned with the ability to generalize our results.
@@
Its compatibility with all transformer-based encoder models makes it easily integrable into various architectures, enhancing its generalization.
@@
VII Related Work
VII-A Pre-trained language models of code
@@
Pre-trained code language models have demonstrated benefits in various SE tasks, including code search, code and documentation generation, and defect prediction [32, 31, 41, 49]. Typically based on the transformer architecture [1], these models can broadly be categorized into decoder-only, encoder-only, and encoder-decoder language models [50].
@@
Decoder-only architectures are typically pre-trained in an autoregressive manner, i.e., predicting the subsequent token based on preceding ones. In the realm of code, examples include InCoder [33], StarCoder 2 [35], and CodeGen 2 [51]. These architectures excel in code generation tasks.
@@
Encoder-only architectures are primarily pre-trained on the masked language modeling task i.e., these models are trained to predict masked words given the remaining sequence as context. Examples include CodeBert [38] and GraphCodeBert [45]. They are frequently used for classification tasks or as a means of feature extraction.
@@
Encoder-decoder architectures undergo pre-training through conditional generation tasks i.e., generating text based on provided input text, e.g., for code summarization or generating code from natural language descriptions. Sample architectures include CodeT5 [32], CodeT5+[52], and plbart [53].
While the experiments in the paper primarily focus on encoder-only models (i.e., CodeBert, GraphCodeBert, and UniXCoder), the proposed pruning layer can be used in both decoder-only and encoder-only architectures. Further experimentation is needed to determine if such layer is also effective in those architectures which is left as future work.
VII-B Optimized transformers
@@
The techniques to enhance the effectiveness of a given transformer model while maintaining its performance fall into three categories—knowledge distillation, quantization, and pruning [20, 28, 54].
@@
Knowledge distillation. The main idea of this technique is to train a smaller model (the student model) to imitate the behavior of a larger model (the teacher model) [55]. In a SE context, Shi et al. [20] introduced Compressor, a framework that employs task-specific knowledge distillation to improve the efficiency of the final transformer model. The authors perform a neural architectural search via a genetic algorithm, followed by training the selected model using knowledge distillation. In their experiments, they use CodeBert [38] and GraphCodeBert [45] as teacher models to tackle the tasks of vulnerability prediction and code clone detection. Compressor still requires the fine-tuning of the teacher model, which incurs overhead throughout the process. Alpine allows to reduce the computational costs of fine-tuning these models, which entails, that such overhead can be reduced.
@@
Quantization. Quantization aims to reduce the memory footprint of neural networks. The core idea is to convert the model’s parameters from 32-bit floating-point numbers (float32 data type) to lower-precision formats, such as 8-bit integers (int8 data type) [56]. Depending on the component of quantization, these techniques can be categorized as quantization of weights only [57, 58, 59] and quantization of both weights and activations [60, 61, 62]. Although quantization reduces the memory footprint of these models, it may also reduce the inference speed and accuracy [63].
@@
Pruning. As mentioned previously, pruning techniques can be divided into weight pruning and token pruning [28]. Liu et al.[64] adopt weight pruning to prune unimportant heads and channels from the mha and ffnn respectively. Zang et al. [22] offer DietCode—a token pruning approach. To build DietCode, Zang et al. carry out an empirical investigation to determine the significance of statements and tokens for CodeBert. Subsequently, based on these findings, they employ a 0-1 Knapsack approach to prune the input sequence. In contrast, our approach is performed adaptively at each layer. Additionally, during the pruning step, we could aggregate the pruned tokens into a single one to prevent excessive loss of information. Finally, our method is language-agnostic, facilitating easy adaptation to diverse programming languages and extending applicability to natural language inputs, particularly beneficial in minimizing computational overhead in tasks like code search or code generation from documentation.
VIII Conclusions and Future Work

### v2

@@ 4. Experiments
Naturally, pruning tokens would result in some information loss. In this research question, we aim to measure the extent of performance drop, in terms of accuracy, for example, that might occur, if any. -
@@
The proposed approach allows one to choose whether pruned tokens are to be partially kept by merging their representation into a single row or entirely removed from the sequence. In this research question, we investigate the impact of such a design choice. -
@@
Finally, an important reason behind model simplification is to allow computationally demanding models to run on relatively lesser-capable gpus (e.g., consumer-grade gpus) compared to those used in high-performance computing clusters (e.g., NVIDIA V or A). This exploration would ensure wider accessibility and make using these models more practical.
4.1. Software Engineering Tasks
In this study, we choose two substantially different tasks from the CodeXGLUE benchmark (Lu et al., 2021) for our experiments. Both tasks aim to improve code quality, developer productivity, software reliability, and maintainability.
Code clone detection: Clone detection identifies similar or identical code fragments, helping maintain code quality, reduce bugs, and enable refactoring. In this work, we used the filtered version of the BigCloneBenchmark dataset (Svajlenko et al., 2014) released by Wang et al. (Wang et al., 2020). This binary classification problem aims to predict whether two Java code snippets are clones. The dataset is composed of training pairs and pairs for testing. We randomly take a subsample of pairs for training and for evaluation; such random sampling is quite common in this domain (Shi et al., 2023; Yang et al., 2022). We use the same subsamples for all experiments during training and evaluation.
Defect prediction: It identifies bugs, vulnerabilities, and defects early, reducing the cost and effort of manual code reviews and testing. We used Zhou et al. (Zhou et al., 2019)’s dataset that was curated from a set of four large-scale open-source C repositories. The authors manually annotated the code snippets and covered multiple types of defects such as resource leaks and use-after-free vulnerabilities. We use the default pre-split sets of samples for training and for validation and testing.
4.2. Model selection
We select the following models given that they represent state-of-the-art encoder-only models in solving the previously mentioned tasks as of conducting this work (Microsoft Research, 2023; Shi et al., 2024; Niu et al., 2023; Zeng et al., 2022b).
CodeBERT (Feng et al., 2020): It is a pre-trained language model that has the same architecture and a similar pretraining strategy as RoBERTa (Liu et al., 2019). It comprises twelve Transformer layers, each containing a 12-headed mha layer and has a hidden dimension of . It was trained on bimodal data consisting of pairs of natural language and programming language across six programming languages from the CodeSearchNet dataset (Husain et al., 2019).
GraphCodeBERT (Guo et al., 2021): GraphCodeBert extends CodeBert by including the data flow information within the source code input. It also includes an additional pre-training task where the objective is to predict whether a data flow edge exists between two nodes.
UniXCoder (Guo et al., 2022): UniXCoder leverages a multi-modal input consisting of natural language, source code, and the flattened sequence of the abstract syntax tree of the code snippet. It is similar to the two other models that share the same internal architecture. The difference here (aside from the pre-training objectives and input representation during pre-training) is that UniXCoder allows for a larger context length, compared to for (Graph)CodeBERT.
@@ 4.2. Model selection
4.3. Evaluation Metrics
To assess the predictive performance of the models on the aforementioned tasks, we use the same evaluation metrics that were reported by Lu et al. (Lu et al., 2021) in the CodeXGLUE benchmark. Specifically, we calculate the accuracy for the Devign dataset and F1-score for the BigCloneBenchmark. As for the computational efficiency, we report the number of floating points operations (flops) and Throughput. As stated in Section 2.2 the number flops quantifies the computational complexity. A model with a higher flop count requires more computing resources. The goal is to reduce such a count while maintaining the predictive performance as much as possible. The throughput refers to the number of input samples a model can process in a second. This metric is especially relevant during inference and model deployment.
4.4. Experimental Setting
@@
5. Results
5.1. RQ1: Computational Efficiency
Impact on FLOPs Count: Table 2(b) highlights the computational speedup offered by Alpine across the three models on the SE tasks. Performing pruning across all layers allows for the most computational savings. On average555Given the values are ratios, we report the geometric mean., it reduces the flops666Calculated using fvcore. count by on the Devign dataset and and BigCloneBenchmark compared the baseline models (i.e., without any pruning). Specifically, CodeBert and GraphCodeBert exhibit roughly the same gain, where these models consume roughly half flops. UniXCoder also show significant gain, though slightly smaller ( and ), in terms of flops.
In Figure 3, we plot the average sequence length across all the twelve layers during inference. When pruning across all layers on the Devign dataset, the sequence length plot exhibits an exponential-like decay on CodeBert and GraphCodeBert. On the other hand, this reduction follows roughly a linear trend on BigCloneBenchmark across all models. What is interesting is that at the final layer, the sequences are quite compressed. Specifically, the highest average sequence length at the last layer () is obtained when using GraphCodeBert on the defect prediction task which is less than the none pruned average length of . In addition, aligned with the results of Table 2(b), we can see that pruning at even-indexed layers yields shorter sequences than odd-indexed ones. Across all task-model combinations and all layers (except for - layer), Alpine produces more compressed sequences when applied at these layers compared to their odd-indexed counterpart.
Impact on Memory Footprint: To further evaluate Alpine’s impact from a computational efficiency standpoint, we report the gpu memory consumption. Due to space constraints, we only report the figures regarding pruning across all layers.
5.2. RQ2: Impact on Performance
In Table 2(b), we report the predictive performance of the pruned models. On the Devign dataset, performing pruning at odd-indexed layers leads to a slight improvement when using CodeBert, with a % increase in accuracy compared to that of the non-pruned version. On the other hand, adding the pruning layer at even layers is more efficient for GraphCodeBert and UniXCoder, where they were able to maintain % and % of the original accuracy scores, respectively. In the clone detection task, pruning at even-indexed layers achieves the highest F1-scores compared to other pruning strategies. It even managed to improve the clone detection performance by increasing the F1-scores of GraphCodeBert and UniXCoder by % and %.
5.3. RQ3: Effect of Merging Prunable Tokens on Model Accuracy
@@
During pruning, the representation of tokens marked to be pruned can be entirely removed or merged into one vector. To study the impact of token merging, we conduct experiments to measure its impact on the two tasks and report the results in Table 3.
@@
5.4. RQ4: Impact of Computing Environment
@@
Building upon these results, we further investigate the implications of the reduced time from an environmental perspective.
In Table 4, we report the CO2 emission rates for both gpus with and without Alpine. The measurements were taken using the tool provided by Lacoste et al. (Lacoste et al., 2019).
@@
6. Discussion
6.1. Sequence Compression
In Section 5.1 we demonstrated how Alpine can significantly reduce sequences’ lengths to less than the original lengths in some scenarios. In code analysis tasks, not all parts of the input sequence contribute equally to the final prediction. Many tokens in source code, such as common language keywords, special characters, or even standard library imports, may not be critical for identifying vulnerabilities or detecting code clones. However, as previously stated in Section 5.1, the reduction rate depends on the nature of the task.
Vulnerability detection, which is the focus of the Devign dataset, usually involves identifying specific code patterns that may lead to security issues. These vulnerability indicators often compromise a small portion of overall function, allowing for rapid pruning of irrelevant code segments. The exponential-like decay observed in Figure 3 suggests that the models may quickly isolate potentially vulnerable code portions, discarding a large amount of non-critical code in earlier layers.
On the other hand, in code clone detection, the task associated with BigCloneBenchmark, requires a more holistic view of code structure and functionality. Clones can manifest in various forms, from exact verbatim copies (i.e., Type-1 clones) to functionality-similar but syntactically different implementations, also known as Type-4 clones. This entails the need to retain a larger set of code features through the layers, which explains the more gradual, linear-like reduction in sequence length. The models likely need to maintain a wider context to effectively compare and identify similarities across different code snippets.
6.2. Less is More
Our findings on sequence reduction with Alpine caters to the calls of developing smaller, task-oriented models (Menzies, 2023). Just as compact models demonstrate that not all parameters in large, general-purpose models are necessary for specific tasks, our results show that not all tokens in an input sequence are crucial for effective code analysis tasks. In both cases, the underlying principle is the same: task-specific efficiency. (Large) language models, while powerful, often contain redundant or task-irrelevant information. Similarly, full code sequences include many tokens that may not contribute significantly to solving SE tasks. This is also similar to previous research. For instance, Xu et al. (Xu et al., 2023b) have observed that removing noisy instances when training models for obsolete comment detection can lead to up to % improvement in accuracy, with % less data. The findings of this work align with the principle of less is more. We can achieve efficient, high-performing systems not by processing more data, but by intelligently selecting the most pertinent information and avoiding noisy data for the task at hand.
6.3. Extension of Alpine to other Transformer Variants
The main focus of Alpine was encoder-only models given their ability to generate embeddings that are primarily suitable for identification tasks. They are also an important component in text generation systems that use information retrieval, or Retrieval Augmented Generation (RAG). However, it is important to remark that Alpine can be extended to other variants of the Transformer architecture, i.e., decoder-only and encoder-decoder models. In decoder-only models, Alpine would be placed after the Masked Self-attention layer (MaskedSA), not after the self-attention layer because the MaskedSA attends only to previous tokens, preserving the autoregressive nature of the decoder. In addition, the last token would be kept, as currently done, so that it can be used by LM head. Finally, in the encoder-decoder setting, Alpine can be either applied to the encoder, decoder, or both. The evaluation of Alpine’s performance on these architectures is deferred to future work. In return, this would increase their adoption given the less computational resources they demand, especially on consumer-grade hardware.
7. Threats to Validity
@@
To mitigate a potential threat, we used the same hyperparameters across all models, as reported in previous works (Lu et al., 2021), and utilized the same data splits for fine-tuning and testing.
@@
Such threats are concerned with the ability to generalize our results.
@@
Its compatibility with all transformer-based encoder models makes it trivial to integrate into various architectures, enhancing its generalization.
@@
8. Related Work
8.1. Pre-trained language models of code
@@
Pre-trained code language models have demonstrated benefits in various SE tasks, including code search, code and documentation generation, and defect prediction (Wang et al., 2021; Guo et al., 2022; Lu et al., 2021; Sharma et al., 2024). Typically based on the transformer architecture (Vaswani et al., 2017), these models can broadly be categorized into decoder-only, encoder-only, and encoder-decoder language models (Min et al., 2023).
@@
Decoder-only architectures are typically pre-trained in an autoregressive manner, i.e., predicting the subsequent token based on preceding ones. In the realm of code, examples include InCoder (Fried et al., 2023), StarCoder 2 (Lozhkov et al., 2024), and CodeGen 2 (Nijkamp et al., 2023). These architectures excel in code generation tasks.
@@
Encoder-only architectures are primarily pre-trained on the masked language modeling task i.e., these models are trained to predict masked words given the remaining sequence as context. Examples include CodeBert (Feng et al., 2020) and GraphCodeBert (Guo et al., 2021). They are frequently used for classification tasks or as a means of feature extraction.
@@
Encoder-decoder architectures undergo pre-training through conditional generation tasks i.e., generating text based on provided input text, e.g., for code summarization or generating code from natural language descriptions. Sample architectures include CodeT5 (Wang et al., 2021), CodeT5+(Wang et al., 2023a), and plbart (Ahmad et al., 2021).
8.2. Optimized transformers
@@
The techniques to enhance the effectiveness of a transformer model while maintaining its performance fall into three categories—knowledge distillation, quantization, and pruning (Shi et al., 2023; Kim et al., 2022; Zhu et al., 2023).
@@
Knowledge distillation. This technique trains a smaller model (the student model) to imitate the behavior of a larger model (the teacher model) (Gou et al., 2021). In a SE context, Shi et al. (Shi et al., 2023) introduced Compressor, a framework that employs task-specific knowledge distillation to improve the efficiency of the final transformer model. The authors perform a neural architectural search via a genetic algorithm, followed by training the selected model using knowledge distillation. In their experiments, they use CodeBert (Feng et al., 2020) and GraphCodeBert (Guo et al., 2021) as teacher models to tackle the tasks of vulnerability prediction and code clone detection. Compressor still requires the fine-tuning of the teacher model, which incurs overhead throughout the process. Alpine allows to reduce the computational costs of fine-tuning these models, minimizing such an overhead.
@@
Quantization. It converts the model’s parameters from 32-bit floating-point numbers (float32 data type) to lower-precision formats, such as 8-bit integers (int8 data type) (Nagel et al., 2021). Depending on the component of quantization, these techniques can be categorized as quantization of weights only (Xu et al., 2023a; Dettmers et al., 2024; Kim et al., 2023) and quantization of both weights and activations (Zafrir et al., 2019; Yao et al., 2022; Dettmers et al., 2023). Although quantization reduces the memory footprint of these models, it may also reduce the inference speed and accuracy (Jin et al., 2024).
@@
Pruning. Pruning techniques can be divided into weight pruning and token pruning (Kim et al., 2022). Liu et al.(Liu et al., 2021) adopt weight pruning to prune unimportant heads and channels from the mha and ffnn respectively. Zang et al. (Zhang et al., 2022) offer DietCode—a token pruning approach. To build DietCode, Zang et al. carry out an empirical investigation to determine the significance of statements and tokens for CodeBert. Subsequently, based on these findings, they employ a 0-1 Knapsack approach to prune the input sequence. In contrast, our approach is performed adaptively at each layer. Additionally, during the pruning step, we could aggregate the pruned tokens into a single one to prevent excessive loss of information. Moreover, our method is language-agnostic, facilitating easy adaptation to diverse programming languages and extending applicability to natural language inputs, particularly beneficial in minimizing computational overhead in tasks like code search or code generation from documentation.
As for LTP (Kim et al., 2022), in addition to the main limitations previously mentioned where concrete computational speed-up is not achieved, there exist some limitations concerning the methodology. First, the reported number of FLOPs is manually calculated, under the assumption that matrix multiplication is optimized when zeroed-out vectors are present. Furthermore, given the placement of the pruning layer, these representations do not remain zeroed-out given the presence of skip connections and normalization layers777This was also pointed out by the issues raised in the replication package repository: https://t.ly/zBMGE and https://t.ly/kLCiU.. Moreover, it does not mask tokens during the evaluation,i.e., the full sequence, threatening the validity of the accuracy results.
9. Conclusions and Future Work